OpenStudy (anonymous):

Solve for x: -3|x + 7| = -12 x = 5 over 3, x = −19 over 3 x = −3, x = −11 x = −3, x = 11 No solution

OpenStudy (igreen):

We break this into two different equations: -3(x + 7) = -12 and -3(x - 7) = -12 ------------------------ -3(x + 7) = -12 Distribute -3 into the parenthesis: -3x - 21 = -12 Add 21 to both sides: -3x = 9 Divide -3 to both sides: x = -3 ------------------------ -3(x - 7) = -12 Distribute -3 into the parenthesis: -3x + 21 = -12 Subtract 21 to both sides: -3x = -33 Divide -3 to both sides: x = -11

OpenStudy (anonymous):

thank u 10000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000 times!!

OpenStudy (zpupster):

This is wrong the answers are right!! Subtract 21 to both sides: -3x = -33 x = -11

OpenStudy (igreen):

No problem.